# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ III.4 Complete system of semi-analytical equations in closed form and their solutions
In Fig. 5 the quasi-particle weight $v_{\mathbf{P}}^{2}$ of the occupied states $\Theta(-E_{\mathbf{P}}^{-})$ is displayed for a number of concentrations corresponding to those in Fig. 4 along with the case $x=1$. In the latter case, the mixture is essentially made of a single component gas of non-interacting point-like fermionic molecules, as illustrated by the Fermi step behavior with occupation number approaching unity. As $x$ decreases, the quasi-particle nature of the occupied states due to the hybridization of the molecular states with atomic ones becomes significant, producing a rather peculiar behavior of the momentum distribution function, as evidenced by its upward bending as the momentum increases, and by the position of the Fermi step, that, in particular for small concentration, is well above the expected position if one would assume the Luttinger theorem to be valid for the composite fermions. This is different with respect to the 3D case, where the Luttinger theorem was found to be respected by the composite fermions in the strong-coupling limit [81]. In that case, however, no hybridization occurred since the condensate vanished above a critical coupling, thus yielding essentially the same picture that we obtain here only at $x=1$. We further notice that in 2D the breakdown of the Luttinger theorem for the hybridized composite fermions is compensated, as far as the integrated quantity $n_{\mathrm{CF}}=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is concerned, by the strong suppression of the weight $v_{\mathbf{P}}^{2}$ when the concentration decreases, in such a way that the identification of $n_{\rm CF}$ with $n_{\rm B}-n_{0}\simeq n_{\rm B}$ is always valid in the strong-coupling limit.
@@
Figure 6: Fermionic momentum distribution as obtained from the strong-coupling limit expression (48) for various bosonic concentrations $x$ at BF coupling $g=2.8$. In the insets the same expression is broken down into the analytical contributions (45), (46a) and (46b), for small and large $x$ where these formulae apply.
Finally, we discuss the profiles of the Fermi momentum distribution (48) for fixed coupling strength and varying concentration. This allows us to confirm the analysis of the molecular and atomic states done above and identify the contributions of paired and unpaired fermions given by expressions (45), (46a) and (46b) of Sec. III.2. In Fig. 6 $n_{\mathrm{F}}({\mathbf{k}})$ is displayed for $g=2.8$ and different values of $x$. For small $x$, a weakly renormalized Fermi step is found in line with a majority of weakly dressed unpaired fermions (46a) and a minority of molecules pictured by the small tail of their internal wave-function (45). As $x$ increases, the height of the Fermi step is reduced by the interaction with the bosons, until $x$ approaches 1 whereby the position of the Fermi step is again associated with the majority fermions, now coming from molecules, but with its height given by the quasi-particle weight $\Delta_{0}^{2}/(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}})^{2}$ of a minority of strongly hybridized unpaired fermions (46b).
The above considerations are quantitatively illustrated in the insets of Fig. 6 for large ($x=1$) and small ($x=0.07$) concentrations. Therein, the distribution function (45) is broken down in paired (CF) and unpaired (UF) contributions, and compared with the exact expression (48). In the top inset we find a small contribution from molecules (CF) and a majority of unpaired fermions (UF), whereas in the bottom panel molecules (CF) are the majority species and unpaired fermions are strongly renormalized (UF).
We thus find a consistent picture throughout the diverse thermodynamic quantities in all regimes originally identified in the different regions of Fig. 3.
IV Numerical results
@@
In this section, we present the fully numerical results for the boson and fermion momentum distributions and for key thermodynamic quantities like the chemical potentials and condensate fraction with varying coupling strength. We also discuss the results for the Tan’s contact parameter [100, 101, 102].
IV.1 Methods
@@
For given values of $n_{\mathrm{F}},n_{\mathrm{B}},a_{\mathrm{BF}}$ and $a_{\rm BB}$ the above system is solved for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$, as follows. First a solution is found for $\mu_{\mathrm{B}}$ from Eq. (53a) with a standard bisection method assuming a suitable ansatz for ${\mu_{\mathrm{F}}}$ and $n_{0}$. Then we consider the 2×2 system of the remaining Eqs. (53b) and (53c) as functions of $\mu_{\mathrm{F}}$ and $n_{0}$ and apply a two-dimensional secant (quasi-Newton) method whereby the approximate Jacobian matrix (which corresponds to the approximate Hessian of the total energy) is updated according to a symmetric rank 1 algorithm [103], that is a generalization of the secant method to multidimensional problems.
An intermediate step of the above procedure is the evaluation of the bosonic and fermionic self-energies, which requires, owing to the low dimensionality of the problem at hand, special care when dealing with the slow convergence of the frequency convolutions appearing in Eqs. (10)-(11). In order to speed up the convergence, the integrand of the convolutions is added and subtracted by an auxiliary function with the same asymptotic behavior yet analytically integrable. As a result, the numerical integration is truncated by applying a large frequency cutoff, which we fix at $\Omega_{c}=\pm 50000{E_{\mathrm{F}}}$ after an accurate analysis of the asymptotic behaviors. For $|\Omega|\geq\Omega_{c}$ integration is done making use of asymptotic expressions whose details are provided in Appendix C.
Figure 7: Bosonic momentum distribution $n_{\mathrm{B}}(k)$ as a function of $k/k_{\mathrm{F}}$, for different values of BF coupling $g$ at boson-boson repulsion $\eta_{\mathrm{B}}=0$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$, (b) $x=0.5$, (c) $x=0.175$. Insets: comparison at large momenta between numerical results and asymptotic behavior (dotted lines) reported in Eq. (61).
@@
Concerning the frequency integral yielding the momentum distributions (14) and (15) we take advantage of the large frequency behavior of the self-energies (78-79) valid in the range $|\omega|\geq 100{E_{\mathrm{F}}}$, whereas for $|\omega|<100{E_{\mathrm{F}}}$ we integrate the full numerical expression (see Appendix C.2). Momentum integration in expressions (16) and (17) is carried out numerically up to a cutoff of $k_{c}=4k_{\mathrm{F}}$, after having subtracted and added a non-interacting or a Bogoliubov Green’s function respectively (see Appendix C.2), and beyond $k_{c}$ by making use of the asymptotic expressions (60-61).
IV.2 Boson and fermion momentum distributions
We now discuss the numerical results for the fermionic and bosonic momentum distributions, Eqs. (14)-(15).
@@ IV.2 Boson and fermion momentum distributions
Figure 8: Bosonic momentum distribution ${n_{\mathrm{B}}}(k)$ divided by the boson concentration $x$ for several values of $x$ and BF coupling $g$.
Figure 9: Bosonic momentum distribution ${n_{\mathrm{B}}}(k)$ evaluated at small momentum $k\simeq 0.003k_{\mathrm{F}}$ as a function of BF coupling $g$, at boson-boson repulsion $\eta_{\mathrm{B}}=0$, for bosonic concentration (a) $x=1$ and (b) $x=0.175$. Triangle: peak value obtained from Eq. (57). Dashed line: guide to the eye.
@@
Figure 7 shows the bosonic momentum distribution from weak to strong BF coupling for different bosonic concentrations $x$ and zero bosonic repulsion, $\eta_{\mathrm{B}}=0$. A peculiar feature common to all panels is the appearance of a peak for small but non-zero momenta, depending on the concentration and coupling. In particular, for very weak coupling the peak persists for all concentrations (full line in all panels) in contrast to higher couplings where it soon disappears. One is thus led to think that a weak-coupling mean-field effect may be at the origin of the peak, primarily due to the unpaired atoms dressing up the weakly interacting bosons, hence a sort of polaronic effect. However, for a definite identification of this peak, an analysis of the bosonic spectral weight function would be in order (this is postponed to future work).
@@
At large $k$, the momentum distribution is expected to be dominated by short-range pairing correlations captured by the asymptotic expression (61) (with the anomalous self-energy turned off in this case). The insets in Fig. 7 compare the numerical large-$k$ behavior of $n_{\mathrm{B}}({\mathbf{k}})$ with the corresponding asymptotic expression (61) (dotted lines).
@@
In strict analogy to what found in 3D [88], and preparing for the universality of the condensate fraction that will be illustrated in Sec. IV.4, it is instructive to inspect the boson momentum distribution divided by the boson concentration $x$. Figure 8 shows that the momentum distributions $n_{\rm B}({\mathbf{k}})$ corresponding to different values of $x$ collapse on top of each other once divided by the concentration $x$, except for a region at small ${\mathbf{k}}$. This universality, which occurs from weak to intermediate coupling ($g\lesssim 0$), suggests that in this coupling range the bosons can be treated as nearly independent of each other (but interacting with the medium), such that $n_{\rm B}({\mathbf{k}})\simeq n_{\rm B}n_{x\to 0}({\mathbf{k}})$, with $n_{x\to 0}({\mathbf{k}})\equiv\lim_{x\to 0}n_{\rm B}({\mathbf{k}})/n_{\rm B}$. We stress that this is not obvious a priori, even in the absence of a direct BB repulsion (as in the case we are considering here), since indirect interactions mediated by the fermions are possible. Indeed, this is what happens at small momenta $k$, as signalled by the deviation of $n_{\rm B}({\mathbf{k}})$ from such independent particle picture in this momentum range.
@@
An interesting feature is also observed when plotting the limiting value for vanishing momentum (${n_{\mathrm{B}}}({\mathbf{k}}\to 0)$) of the momentum distribution of non-condensed bosons as a function of coupling, as shown in Fig. 9 for two representative values at large and small boson concentrations. By keeping $k/k_{\mathrm{F}}$ fixed at small value and scanning the BF interaction, a prominent peak appears close to $g\simeq-0.7$ which eventually diverges in the limit $k=0$. We stress that in the present case the BB repulsion is turned off and thus there is no anomalous self-energy term (which would yield $n_{\rm B}({\mathbf{k}}\to 0)\to\infty$ for all values of $g$, as standard from Bogoliubov theory).
@@
$\Sigma_{\mathrm{B}}({\mathbf{k}},\omega)=\Sigma_{\rm B}(0,0)+a\,i\omega+b\,\omega^{2}+c\,k^{2},$ (54)
@@
where $1/m_{\mathrm{B}}^{*}=1/m_{\mathrm{B}}+c$ and the Hugenholtz-Pines condition has been used. By fitting the coefficients $a,b,c$ with the data for the boson self-energy at small $\omega$ and ${\mathbf{k}}$, we find that, at $g\simeq-0.7$, $a=1$ and $b$ is positive, resulting in the following expression for the momentum distribution (15)
$n_{\mathrm{B}}({\mathbf{k}})\simeq\int_{-\omega_{\rm IR}}^{+\omega_{\mathrm{IR}}}\frac{d\omega}{2\pi}\frac{1}{{{\mathbf{k}}^{2}\over 2m_{\mathrm{B}}^{*}}+b\omega^{2}}\\ +2\int_{\omega_{\mathrm{IR}}}^{+\infty}\frac{d\omega}{2\pi}\Re[G_{\mathrm{B}}({\mathbf{k}},\omega)],$ (56)
@@
Figure 10: Bosonic momentum distribution $n_{\mathrm{B}}(k)$ as a function of momentum $k/k_{\mathrm{F}}$, for different values of boson-fermion coupling $g$ at boson-boson repulsion $\eta_{\mathrm{B}}=0.1$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$; (b) $x=0.5$; (c) $x=0.175$.
@@
Figure 11: Fermionic momentum distribution for bosonic concentrations (a) $x=1$, (b) $x=0.5$ and (c) $x=0.175$ and different values of boson-fermion coupling at boson-boson repulsion $\eta_{\mathrm{B}}=0$. Symbols indicate numerical data, while lines are their analytical benchmark. The same thermodynamical parameters are employed in the calculation of both sets of curves. Insets: same quantities in log-scale.
@@
For this reason one needs to include a BB repulsion in order to stabilize the system against the aforementioned effects. We thus consider the case where a direct BB repulsion is turned on, causing a divergence at zero $k$ to be present at all $g$’s, as expected within Bogoliubov theory. Calculations for a relatively strong boson-boson repulsion $\eta_{\mathrm{B}}=0.1$ are reported in Fig. 10. Clearly, the main differences with respect to the case of zero repulsion occur at small momenta, where the occupation of low momenta states predicted by Bogoliubov theory dominates. Regarding the large momentum regime (not reported in Fig. 10), it is well described by expression (61) discussed below.
@@
Concerning the Fermi momentum distribution we first discuss its comparison with the strong-coupling expression (48) with a two-fold aim: to untangle the different contributions from bare or dressed atoms and molecules based on the lesson learned in Sec. III and, at the same time, to perform a stringent check on our numerical implementation. Fig. 11 displays the numerical results for ${n_{\mathrm{F}}}({\mathbf{k}})$ as symbols and the expression (48) as lines for several values of BF coupling $g$ and (a) $x=1$, (b) $x=0.5$ (c) $x=0.175$, using the same thermodynamic parameters for both sets of curves. The agreement is remarkably good for $g\gtrsim 0$, thus demonstrating that the fermionic Green’s function is well captured by the BCS-like expression (32) in an extended coupling range and not only for large values of $g$.
Figure 12: Fermionic momentum distribution ${n_{\mathrm{F}}}(k)$ as a function of momentum $k/k_{\mathrm{F}}$, for different values of BF coupling $g$ at boson-boson repulsion $\eta_{\mathrm{B}}=0$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$, (b) $x=0.5$, (c) $x=0.175$. Insets: ratio of ${n_{\mathrm{F}}}(k)$ to its asymptotic limit given by Eq. (60).
@@
We note that the discussion on paired and unpaired fermions of Sec. III applies here, with the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$ representing the molecular wave-function and the Fermi step measuring the amount of unpaired yet correlated fermions. We notice that, when $g$ is reduced, discrepancies between the asymptotic curve and numerical data occur first for higher boson concentration, and they occur in the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$. This is related to the onset of the continuum of particle-hole excitations becoming comparable to the molecular binding energy, thus making the BCS-like expression (48) less valid.

### v2

@@ 3.4 Complete system of semi-analytic equations in closed form and their solutions
The momentum distribution of the composite fermions $n_{\mathrm{CF}}(\mathrm{P})=v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is displayed in Fig. 5 for a number of concentrations corresponding to those in Fig. 4 along with the case $x=1$. In the latter case, the mixture is essentially made of a single component gas of non-interacting point-like fermionic molecules, as illustrated by the Fermi step behavior with occupation number approaching unity. As $x$ decreases, interaction effects, originating from the hybridization of the molecular and atomic states, become significant and generate a rather peculiar behavior of the momentum distribution function, as evidenced by its upward bending with increasing momentum. (A similar behavior was also found in [71] for a BF mixture in a 1D lattice.) Moreover, we notice that the Fermi step, in particular for small concentrations, is located well above the expected position, if one would assume the validity of the Luttinger theorem for the composite fermions.
The Luttinger theorem states that the volume of the Fermi sphere of a Fermi liquid does not depend on interaction, and thus coincides with the volume of the Fermi sphere of the corresponding non-interacting Fermi gas [113]. Its validity has been extended to density-imbalanced two-component Fermi gases [114] and, within a two-channel model, to BF mixtures [65]. In both cases, it was shown that, provided the system is in the normal phase, the volumes of the Fermi spheres are separately conserved for the two Fermi components (in the BF mixture case, the two components correspond to renormalized fermionic atoms and molecules). This is indeed what is found in the strong-coupling limit of a 3D BF mixture in [84], as in this limit the condensate vanishes identically. In contrast, in the present 2D case the condensate density never identically vanishes, and in particular the hybridization energy $\Delta_{0}$ remains finite even for $g\to\infty$ (except in the trivial case $x=0$ and in the more interesting case $x=1$).
In a BF mixture with a condensate, it was shown in [65] that it is the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules that should be conserved. Fig. 5 shows a clear violation of the latter version of the Luttinger theorem since the molecular Fermi momentum is always larger than $k_{\mathrm{F}}$, which is the Fermi momentum that should determine the total volume (i.e the sum of the volumes of the Fermi spheres of dressed atomic fermions and molecules). We ascribe this shortcoming to two factors. First, the large values of the hybridization energy-scale $\Delta_{0}$ obtained in our theory, which are in turn caused by an overestimation of the atom-molecule repulsion (see Sec. 3.1), eventually lead to an overpopulation of the condensate. So, the simple strong-coupling picture of a Fermi-Fermi mixture of unpaired fermions and molecules obeying the Luttinger theorem is never reached in 2D. Second, as stressed in [114] and explicitly shown in [115], when dealing with approximate theories the validity of the Luttinger theorem is not guaranteed for schemes, like the present one, that do not use fully self-consistent fermionic Green’s functions in their formulation.
@@
Figure 6: Fermionic momentum distribution $n_{\rm F}({\mathbf{k}})$ as a function of $k/k_{\rm F}$, as obtained from the strong-coupling asymptotic expression (39), for various bosonic concentrations $x$, at BF coupling $g=2.8$ and BB repulsion $\eta_{\mathrm{B}}=0$. In the insets, the same expression (full line) is compared with the analytical contributions obtained in the limit of small hybridization: $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$ (CF) and $n_{\rm UF}({\mathbf{k}})$ (UF) as given by Eq. (43), and their sum (tot).
Nevertheless, the breakdown of the Luttinger theorem in 2D for the hybridized composite fermions is compensated, as far as the integrated quantity $n_{\mathrm{CF}}=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}n_{\rm CF}({\mathbf{P}})=\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})$ is concerned, by the strong suppression of the weight $v_{\mathbf{P}}^{2}$ when the concentration decreases, in such a way that the identification of $n_{\rm CF}$ with $n_{\rm B}-n_{0}\simeq n_{\rm B}$ is always valid in the strong-coupling limit.
Finally, we discuss the profiles of the Fermi momentum distribution (39) for fixed coupling strength and different concentrations. In Fig. 6, $n_{\mathrm{F}}({\mathbf{k}})$ is displayed for $g=2.8$ and different values of $x$. For small $x$, most fermions remain unpaired and are described by a weakly renormalized Fermi step, while a small fraction of fermions bound in molecules are described by the small contribution $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$, with $n_{\rm CF}\ll n_{\rm F}$. As $x$ increases, the height of the Fermi step is progressively reduced, indicating the increasing importance of interaction effects. The position of the Fermi step, on the other hand, decreases only slightly from the non-interacting value $k_{\rm F}$. We notice that this behavior, together with the one just discussed for $n_{\rm CF}({\mathbf{P}})$ reflects again the breakdown of the Luttinger theorem that, in its modified version in the presence of a condensate [65], would require $k_{\rm F}^{2}=k_{\rm F,UF}^{2}+P_{\rm F,CF}^{2}$ (where $k_{\rm F,UF}$ and $P_{\rm F,CF}$ indicate the Fermi momenta for the unpaired and composite fermions, respectively).
An important exception is the case $x=1$. In this case, the height of the Fermi step for $n_{\rm F}({\mathbf{k}})$ is given by the quasi-particle weight $\Delta_{0}^{2}/(\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}-\tilde{\xi}_{\mathbf{k}})^{2}$ (cf. Eq. (D.22b) of Appendix D.2), with $\Delta_{0}$ vanishing in the limit $g\to\infty$ (see also Fig. 14 (b) below). The Fermi step thus vanishes, effectively corresponding to $k_{\rm F,UF}=0$, and $n_{\rm F}({\mathbf{k}})$ is entirely described by the contribution $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$ from fermions bound into molecules. At the same time, $P_{\rm F,CF}=k_{\rm F}$, as one can see in Fig. 5, consistently with the Luttinger theorem for a system made only by molecules.
The insets of Fig. 6 finally compare the full strong-coupling expression (39) for $n_{\rm F}({\mathbf{k}})$ with its small $\Delta_{0}/E_{\rm F}$ expansion, Eqs.(41-43), which holds for $x\to 0$ and $x\to 1$, as mentioned above. The distribution function (41) resulting from this expansion (tot) is broken down in paired (CF) and unpaired (UF) contributions, described by $n_{\rm CF}|\phi({\mathbf{k}})|^{2}$ and Eq. (43) for $n_{\rm UF}({\mathbf{k}})$, respectively. One sees from this comparison that the small $\Delta_{0}/E_{\rm F}$ expansion fully agrees with the full strong-coupling expression (39). One also sees how, for $x=1$, the unpaired contribution is already strongly suppressed for the value $g=2.8$ considered in Fig. 6, consistently with its vanishing contribution for $g\to\infty$ discussed above.
4 Numerical results
@@
In this section, we present the fully numerical results for the boson and fermion momentum distributions and for key thermodynamic quantities like the chemical potentials and condensate fraction with varying coupling strength. We also discuss the results for Tan’s contact parameter [116, 117, 118]. In all calculations we restrict to the case of equal boson and fermion masses: $m_{\rm B}=m_{\rm F}$.
4.1 Methods
@@
For given values of $n_{\mathrm{F}},n_{\mathrm{B}},a_{\mathrm{BF}}$ and $a_{\rm BB}$ the above system is solved for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$, as follows. First, a solution is found for $\mu_{\mathrm{B}}$ from Eq. (48a) with a standard bisection method assuming a suitable ansatz for ${\mu_{\mathrm{F}}}$ and $n_{0}$. Then we consider the 2×2 system of the remaining Eqs. (48b) and (48c) as functions of $\mu_{\mathrm{F}}$ and $n_{0}$ and apply a two-dimensional secant (quasi-Newton) method whereby the approximate Jacobian matrix (which corresponds to the approximate Hessian of the total energy) is updated according to a symmetric rank 1 algorithm [119], that is a generalization of the secant method to multidimensional problems.
An intermediate step of the above procedure is the evaluation of the bosonic and fermionic self-energies, which requires, owing to the low dimensionality of the problem at hand, special care when dealing with the slow convergence of the frequency convolutions appearing in Eqs. (11)-(12). In order to speed up the convergence, the integrand of the convolutions is added and subtracted by an auxiliary function with the same asymptotic behavior yet analytically integrable. As a result, the numerical integration is truncated by applying a large frequency cutoff, which we fix at $\Omega_{c}=\pm 50000{E_{\mathrm{F}}}$ after an accurate analysis of the asymptotic behaviors. For $|\Omega|\geq\Omega_{c}$ integration is done by making use of asymptotic expressions whose details are provided in Appendix C.
Figure 7: Bosonic momentum distribution $n_{\mathrm{B}}(k)$ as a function of $k/k_{\mathrm{F}}$, for different values of BF coupling $g$ and BB repulsion $\eta_{\mathrm{B}}=0$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$, (b) $x=0.5$, (c) $x=0.175$. Insets: comparison between numerical results and large momentum behavior described by Eq. (49) (dotted lines).
@@
Concerning the frequency integral yielding the momentum distributions (15) and (16) we take advantage of the large frequency behavior of the self-energies (C.5-C.6) valid in the range $|\omega|\geq 100{E_{\mathrm{F}}}$, whereas for $|\omega|<100{E_{\mathrm{F}}}$ we integrate the full numerical expression (see Appendix C.2). Momentum integration in expressions (17) and (18) is carried out numerically up to a cutoff of $k_{c}=4k_{\mathrm{F}}$, after having subtracted and added a non-interacting or a Bogoliubov Green’s function respectively (see Appendix C.2), and beyond $k_{c}$ by making use of the asymptotic expressions (56-57).
@@
Figure 8: Bosonic momentum distribution ${n_{\mathrm{B}}}(k)$ divided by the boson concentration $x$ as a function of $k/k_{\mathrm{F}}$, for several values of $x$ and BF coupling $g$.
4.2 Boson and fermion momentum distributions
@@
We now discuss the numerical results for the fermionic and bosonic momentum distributions, Eqs. (15)-(16). Figure 7 shows the bosonic momentum distribution from weak to strong BF coupling for different bosonic concentrations $x$ and zero bosonic repulsion, $\eta_{\mathrm{B}}=0$. A peculiar feature common to all panels is the appearance of a peak for small but non-zero momenta, depending on the concentration and coupling. In particular, for very weak coupling the peak persists for all concentrations (full line in all panels) in contrast to higher couplings where it soon disappears. One is thus led to think that a weak-coupling mean-field effect may be at the origin of the peak, primarily due to the unpaired atoms dressing up the weakly interacting bosons, hence a sort of polaronic effect. However, for a definite identification of this peak, an analysis of the bosonic spectral weight function would be in order (this is postponed to future work).
@@
At large $k$, the momentum distribution is expected to be dominated by short-range pairing correlations captured by the asymptotic expression (see Sec. 4.5)
${n_{\mathrm{B}}}(k)\simeq\frac{\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}T({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}}{\left(\frac{k^{2}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)^{2}}.$ (49)
The insets in Fig. 7 compare the numerical large-$k$ behavior of $n_{\mathrm{B}}({\mathbf{k}})$ with the corresponding asymptotic expression (49) (dotted lines).
Figure 9: Bosonic momentum distribution ${n_{\mathrm{B}}}(k)$ evaluated at small momentum $k\simeq 0.003k_{\mathrm{F}}$ as a function of BF coupling $g$, at BB repulsion $\eta_{\mathrm{B}}=0$, for bosonic concentration (a) $x=1$ and (b) $x=0.175$. Triangle: peak value obtained from Eq. (53). Dashed line: guide to the eye. Insets: parameters $a,b,c$ of Eq. (50), where the dotted lines are linear fits.
@@
In strict analogy to what is found in 3D [91], and preparing for the universality of the condensate fraction that will be illustrated in Sec. 4.4, it is instructive to inspect the boson momentum distribution divided by the boson concentration $x$. Figure 8 shows that the momentum distributions $n_{\rm B}({\mathbf{k}})$ corresponding to different values of $x$ collapse on top of each other once divided by the concentration $x$, except for a region of small ${\mathbf{k}}$. This universality, which occurs from weak to intermediate coupling ($g\lesssim 0$), suggests that in this coupling range the bosons can be treated as nearly independent of each other (but interacting with the medium), such that $n_{\rm B}({\mathbf{k}})\simeq n_{\rm B}n_{x\to 0}({\mathbf{k}})$, with $n_{x\to 0}({\mathbf{k}})\equiv\lim_{x\to 0}n_{\rm B}({\mathbf{k}})/n_{\rm B}$. We stress that this is not obvious a priori, even in the absence of a direct BB repulsion (as in the case we are considering here), since indirect interactions mediated by the fermions are possible. Indeed, this is what happens at small momenta $k$, as signaled by the deviation of $n_{\rm B}({\mathbf{k}})$ from such an independent particle picture in this momentum range.
Figure 10: Bosonic momentum distribution $n_{\mathrm{B}}(k)$ as a function of $k/k_{\mathrm{F}}$, for different values of boson-fermion coupling $g$ at BB repulsion $\eta_{\mathrm{B}}=0.1$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$; (b) $x=0.5$; (c) $x=0.175$.
@@
An interesting feature is also observed when plotting the limiting value for vanishing momentum (${n_{\mathrm{B}}}({\mathbf{k}}\to 0)$) of the momentum distribution of non-condensed bosons as a function of coupling, as shown in Fig. 9 for two representative values at large and small boson concentrations. By keeping $k/k_{\mathrm{F}}$ fixed at a small value and scanning the BF interaction, a prominent peak appears close to $g\simeq-0.7$ which eventually diverges in the limit $k=0$. We stress that in the present case, the BB repulsion is turned off and thus there is no anomalous self-energy term (which would yield $n_{\rm B}({\mathbf{k}}\to 0)\to\infty$ for all values of $g$, as standard from Bogoliubov theory).
@@
$\Sigma_{\mathrm{B}}({\mathbf{k}},\omega)=\Sigma_{\rm B}(\textbf{0},0)+a\,i\omega+b\,\omega^{2}+c\,k^{2},$ (50)
@@
where $1/m_{\mathrm{B}}^{*}=1/m_{\mathrm{B}}+c$ and the Hugenholtz-Pines condition has been used. The coefficients $a,b,c$ are obtained by fitting the data for the bosonic self-energy at small $\omega$ and ${\mathbf{k}}$, for fixed values of $g$. Then, through a simple linear regression, one can obtain $a,b,c$ as a function of $g$, as shown in the insets of Fig. 9. One can notice that for $g\simeq-0.7$, $a=1$ and $b>0$, resulting in the following expression for the momentum distribution (16)
$n_{\mathrm{B}}({\mathbf{k}})\simeq\int_{-\omega_{\rm IR}}^{+\omega_{\mathrm{IR}}}\!\frac{d\omega}{2\pi}\frac{1}{{{\mathbf{k}}^{2}\over 2m_{\mathrm{B}}^{*}}+b\omega^{2}}+2\int_{\omega_{\mathrm{IR}}}^{+\infty}\!\frac{d\omega}{2\pi}\Re[G_{\mathrm{B}}({\mathbf{k}},\omega)],$ (52)
@@
Figure 11: Fermionic momentum distribution $n_{\rm F}(k)$ as a function of $k/k_{\mathrm{F}}$ for bosonic concentrations (a) $x=1$, (b) $x=0.5$ and (c) $x=0.175$, for different strong-coupling values of the BF coupling $g$ and BB repulsion $\eta_{\mathrm{B}}=0$. Symbols indicate numerical data, while lines correspond to the strong-coupling approximation (39) using the same thermodynamical parameters as in the numerical calculation. Insets: same quantities in log-scale.
@@
For this reason, one needs to include a BB repulsion in order to stabilize the system against the aforementioned effects. We thus consider the case where a direct BB repulsion is turned on, causing a divergence at zero $k$ to be present at all $g$’s, as expected within Bogoliubov theory. Calculations for a relatively strong boson-boson repulsion $\eta_{\mathrm{B}}=0.1$ are reported in Fig. 10. Clearly, the main differences with respect to the case of zero repulsion occur at small momenta, where the occupation of low momenta states predicted by Bogoliubov theory dominates. Regarding the large momentum regime (not reported in Fig. 10), it is well described by expression (57) discussed below.
@@
Concerning the Fermi momentum distribution, we first discuss its comparison with the strong-coupling expression (39) with a two-fold aim: to untangle the different contributions from atomic and molecular fermions based on what we learned in Sec. 3.2 and, at the same time, to check to what extent the approximate expression (38) for $G_{\rm F}({\mathbf{k}},\omega)$, from which Eq. (39) for ${n_{\mathrm{F}}}({\mathbf{k}})$ immediately results, provides a good approximation for $G_{\rm F}({\mathbf{k}},\omega)$. Fig. 11 displays the numerical results for ${n_{\mathrm{F}}}({\mathbf{k}})$ as symbols and the expression (39) as lines for several values of BF coupling $g$ and (a) $x=1$, (b) $x=0.5$ (c) $x=0.175$, using the same thermodynamic parameters for both calculations. The use of the same thermodynamic parameters allows for a separate check of the approximation (38) for $G_{\rm F}({\mathbf{k}},\omega)$, independently of the remaining strong-coupling approximations (46) and (47) affecting the values of the thermodynamic parameters. The agreement is remarkably good for $g\gtrsim 1$, and remains reasonable even at the lowest coupling $g=0.4$ considered in Fig. 11. This demonstrates that the fermionic Green’s function is well captured by the “BCS-like” expression (38) in an extended coupling range and not only for large values of $g$, where it is expected to hold.
@@
We also note that the discussion on paired and unpaired fermions of Secs. 3.2 and 3.4 applies here, with the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$ representing the internal part of the molecular wave-function and the Fermi step being associated with unpaired yet correlated fermions. We notice that, when $g$ is reduced, discrepancies between the asymptotic curve and numerical data occur first for higher boson concentration, and they occur in the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$. This is related to the onset of the continuum of particle-hole excitations becoming comparable to the molecular binding energy, thus making the BCS-like expression (39) less valid.
Figure 12: Fermionic momentum distribution ${n_{\mathrm{F}}}(k)$ as a function of $k/k_{\mathrm{F}}$, for different values of the BF coupling $g$ and BB repulsion $\eta_{\mathrm{B}}=0$. Different cases of bosonic concentration $x$ are reported: (a) $x=1$, (b) $x=0.5$, (c) $x=0.175$. Insets: ratio of ${n_{\mathrm{F}}}(k)$ to its large-momentum behavior given by Eq. (56).